OpenStudy (owlcoffee):

What is xyz if: \[x^2 y z^3 =w^3\] \[xy^2 = w^9 \]

TheSmartOne (thesmartone):

\(x^2yz^3 \times xy^3 = x^3y^3w^3 = w^3 \times w^9 = w^{12}\) \(x^3y^3z^3 = w^{12}\) \(\sqrt[3]{x^3y^3z^3} = \sqrt[3]{w^{12}}\) \(xyz = w^4\)